Session: Nov 1998 (session 4)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Orfei at Sun Nov 15 01:43:40 MET 1998) Success
LCP channel available all time. RCP channel available from 13:25UT. No known other problems.

Medicina (A. Polatidis at Sun Nov 15 12:09:02 MET 1998) Success
No known problems

Torun (K.Borkowski at Mon Nov 16 14:12:47 MET 1998) Minor failures
9:36 to 9:43 UT lost due to tape movement problem. 12:38 - early stop of the tape by the low-tape sensor (about 2 minutes of the scan lost).

Effelsberg (dgraham at Mon Nov 16 16:46:06 MET 1998) Minor failures
elevation too low at start of ON2. Stopped for snow 1344-1357

Jodrell (Alastair Gunn at Wed Nov 18 11:27:19 MET 1998) Success
No known problems.

Calibration (Jean\-Francois Desmurs at Tue Nov 24 19:07:26 MET 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the nov98 directory under the name: ep023.antab.Read the comments in the file to get further information.